OpenStudy (anonymous):

Part A: Solve F = 9/5= C + 32 for C. Part B: Determine the value of C when F = 10 °F. Part C: Solve -np - 70 > 40 for n. Show your work

OpenStudy (anonymous):

Part A:9/5c+32=F -32 to both sides 9/5c=F-32 this is as far as I got in part A I don't know what I should do next

OpenStudy (anonymous):

multiply both sides by \(\frac{5}{9}\)

OpenStudy (anonymous):

you can even leave the parentheses there and write \[C=\frac{5}{9}(F-32)\]

OpenStudy (anonymous):

O okay cool, thanks! Now can you help me with part B and C?

OpenStudy (anonymous):

now that you have \[C=\frac{5}{9}(F-32)\] if \(F=10\) you can compute \[C=\frac{5}{9}(10-32)\]

OpenStudy (anonymous):

you okay with that?

OpenStudy (anonymous):

Yes Okay but for part C is it even really related to part A and B at all

OpenStudy (anonymous):

no it isn't but we can do that as well

OpenStudy (anonymous):

oh actually we can't!

OpenStudy (anonymous):

we can start with \[ -np - 70 > 40\] add \(70\) and get \[-np>110\]but after that we cannot solve for \(n\)

OpenStudy (anonymous):

the reason you cannot solve for \(n\) is that you do not know if \(p\) is positive or negative you want to divide by \(-p\) but you cannot, because if \(-p>0\) you leave the inequality alone, but if \(-p<0\) you have to switch the inequality

OpenStudy (anonymous):

O okay that makes sense....thank you bunches

OpenStudy (anonymous):

welcome bunches